# Real-Time Image Segmentation via Hybrid Convolutional-Transformer Architecture Search

### v1

@@ III-C Searching with Multi-objective
Inspire by [40] and [41], we formulate the searching of the optimal architecture $\alpha^{*}$ as a two-stage optimization problem. The architecture search space $\mathcal{A}$ is encoded in the supernet, denoted as $\mathcal{N}(\mathcal{A},\mathcal{W})$. Each candidate architectures $\alpha\in\mathcal{A}$ share their weights across $\mathcal{W}$. In the first stage, we optimize the supernet weights $\mathcal{W}$ by:
$\mathcal{W}=\mathop{arg\min}\limits_{\mathcal{W}}\mathcal{L}_{train}(\mathcal{N}(\mathcal{A},\mathcal{W}))$ (1)
@@
The prior distribution of $\Gamma(\mathcal{A})$ is important. In order to meet our goal of searching for architectures with arbitrary branch from different resolution, we pre-define this prior in the training process of the supernet. Since the sub-network with more branches contains the sub-network with less branches, we empirically set the sub-network with more branches to have a higher sampling probability.
$\rho_{i}<\rho_{i+1},i\in{1,2,3}$ (2)
where $\rho_{i}$ means the sampling of probability of architecture contains $i$ branches. Other components in sub-network $\alpha$ are sampling from uniform distribution.
Once the supernet training is finished, the second-stage is to search for architectures via ranking the accuracy and latency of each sub-network $\alpha\in\mathcal{A}$. We convert the architecture searching problem to a standard multi-objective optimization (MOP) problem.
@@
Where $\alpha$ represents the decision vector, i.e., the searched network architecture, $y$ represents the target vector, i.e., the mIoU and latency. $\mathcal{A}$ represents the search space formed by the decision vector $\alpha$, and $Y$ represents the target space formed by the target vector $y$. And the constraint $e(x)\leq 0$ determines the feasible value range of the decision vector. In order to ensure that the searched architectures are effective, we set the following three constraints: 1) The number of paths in each sub-network needs to be greater than 0; 2) For a sub-network that contains multiple paths, the locations of down-sampling should be different; 3) In the down-sampling position, skip-connection cannot be selected; 4) The memory usage of a sub-network during training cannot exceed hardware limit.
Different from the Single-objective Optimization Problem, there is usually no unique optimal solution in MOP problem, but a Pareto optimal solution set. Hence, we adapt the popular evolution algorithm NSGA-II [42] to solve the optimization problem. The details of our searching algorithm can be found in the supplemental material.
The detail searching algorithm are presented in Algorithm  1. One-hot encoding is employed to encode the sub-network $\alpha$. When the sampling is over, we will decode the sub-network to evaluate its corresponding mIoU and latency. We solve this multi-objective optimization (MOP) problem under the NSGA-II [42] framework.
@@
IV Experiments
We evaluate our method on the semantic segmentation task.
Figure 3: An illustration of our searched architectures with multiple branches. The red dot indicates the location of the self attention module.
IV-A Benchmark and Evaluation Metrics
@@
Cityscapes [43] is a popular dataset which contains a diverse set of stereo video sequences recorded in street scenes from 50 different cities. It has 5,000 high quality pixel-level annotations. There are 2,975 images for training, 500 images for validation. And for testing, it offers 1,525 images without ground-truth for a fair comparison. The dense annotation contains 19 classes for each pixel.
@@
Evaluation Metrics. The mean intersection over union per class (mIoU) and FPS (frame per second) are used as the metrics for semantic segmentation and speed, respectively.
IV-B Setup
@@
Search. During training of the supernet, the sampling ratios for $\rho_{1}$, $\rho_{2}$ and $\rho_{3}$ is set to 0.2, 0.3 and 0.5, respectively. The initial channel for supernet is 64 and the mini-batch size is set to 16. The supernet is trained on the Cityscapes dataset with $256\times 512$ images on the training set. We use SGD optimizer with momentum 0.9 and weight decay of $5\times 10^{-4}$ to optimize the supernet weights. The learning rate is set to 0.02 with the exponential learning rate decay of 0.992. The entire supernet optimization takes about 120 hours on a single V100 GPU. In the search process, The initial population $P_{0}$ is set to 40 and generated by uniform random sampling. The probabilities of crossover and mutation operations are set at 0.9 and 0.02 respectively. Before the inference of a sub-network, the statistics of all the Batch Normalization [44] operations are recalculated on a random subset of training data (1000 images in Cityscapes). The number of generations is 20 and we search sub-networks with different branch numbers separately. Hence, a total of 2400 network sub-networks are searched.
@@
Retrain. Similar to [14], we also use the filter multiplier to control the model capacity, and the default value for HyCTAS-S is set as 12, respectively. The final output stride of HyCTAS-S is 8 by merging the low level features in stem. We use SGD optimizer with momentum 0.9 and weight decay of $5\times 10^{-4}$ to optimize the searched networks. The ’poly’ learning rate policy is employed with an initial learning rate of 0.005. Data augmentations on all datasets, i.e., random crop, flipping etc., are consistent with [45]. On the Cityscapes dataset, the training images are cropped to $512\times 1024$ and the mini batch size is set to 32. It is worth noting that all our results are trained from scratch without ImageNet pretraining and we only use the training set without any extra data.
@@
Evaluation. We use Nvidia Geforce RTX 3090 and Intel Xeon E5-2680-v4 CPU (2.40GHz) for benchmarking the computation cost. When testing the inference speed, an image of a batch size of 1 is first loaded into the GPU memory and the model is warmed up to reach a steady speed. After that, the inference time is measured by running the model for six seconds. All the reported results, unless specified, are evaluated without multi-scale inference and flipping. The mean intersection over union per class (mIoU) is used as the evaluation metric for semantic segmentation.
@@
TABLE III: Comparison of semantic segmentation on the Cityscapes. $\dagger$: speed on NVIDIA Xavier, $\sharp$: speed with TensorRT acceleration, $\star$: trained from scratch.
@@
IV-C Model Exploration
Our exploration focuses on balancing ”Speed v.s. mIoU” in architecture search, highlighting how our framework excels in optimizing multiple objectives. Figure 3 shows architecture evolution, with color-coded branch numbers and a highlighted Pareto boundary at the search’s end, demonstrating the effect of dual-goal targeting on architecture distribution. We prioritize architectures on the Pareto frontier for their optimal balance and practical relevance.
We visualized searched architectures with varying branch numbers in Figure 3 (a)(b)(c), revealing interesting patterns in the placement of self-attention modules. With a single branch, self-attention is positioned at higher resolutions to enhance performance. Conversely, with three branches, self-attention is placed deeper to effectively fuse features across branches, aligning well with its strengths and limitations.
IV-D Comparison to SOTA
@@
Cityscapes. Shown in Table III, we can see that when the speed is more than 100FPS, our model establishes new state-of-the-arts. HyCTAS-M consistently outperforms recent BiSeNet [25], Fast-SCNN [46] and DF1-Seg-d8 [47]. In particular, our HyCTAS-M achieves 77.1% mIoU on Cityscapes validation set with 29.4G Flops and 2.4M parameters, which surpasses HRNetV2-W18-Small-v1 by 6.8% with similar model size. On the Cityscapes test set, our HyCTAS-M surpasses FasterSeg [15] by 4.2 %. Moreover, the speed of our model reaches an astonishing 87.8 FPS without TensorRT acceleration. Meanwhile, our HyCTAS-S has the smallest Flops and parameters and achieves the fastest inference speed at 159.2 FPS. This proves that our models are very competitive when deployed on real hardware.
Our method outperforms recent NAS models in search efficiency and GPU memory usage. Firstly, our multi-objective strategy allows HyCTAS to find models catering to varied needs, such as low latency or high performance, within just 5 GPU days. In contrast, traditional NAS methods search for a single model at a time, leading to linearly increasing search costs with each additional model. Secondly, HyCTAS-M surpasses baseline NAS models in performance; for instance, it’s 3.8% better than FasterSeg [15] on the Cityscapes validation set.
These results are achieved with only Cityscapes fine-annotated images, without using any extra data (coarse-annotated images, COCO, ImageNet, etc.).
IV-E Ablation
@@
Compared to Random Sampling. To evaluate our search strategy, we contrasted our architectures with 5 randomly sampled ones (1-3 branches) on the Cityscapes validation set. Table IV shows the average performance of these random architectures. Our discovered Searched-1 architecture surpasses its random counterpart by 4.3% in accuracy with a reduced model size. The discovered Searched-2 and Searched-3 both surpass random architectures by more than 3%.
@@
Robustness of Searching. We further conduct an ablation study by intentionally changing components to verify the searching robustness on our HyCTAS-M model. Shown in Table V, we distort the searched architectures by replacing all self-attention to convolution, replacing all convolution to self-attention, shifting self-attention to another position and shifting self-attention to another branch, i.e. moved it to a random position on the same stride. It can be seen that our searched models consistently perform better these variants and proves its robustness.
@@
TABLE IV: Ablation study on the random sampled
architectures on the Cityscapes validation set.
@@
TABLE V: Ablation study on the searched architectures on the Cityscapes validation set.
@@
TABLE VI: Experiments with different settings on Cityscapes validation set. ‘SDP’ denotes Scheduled Drop Path.
@@
Retrain Settings. From Table  VI, we can see that adding the scheduled drop path [51, 14] improves the mIoU by 1.5%. When set the scheduled drop path probability to 0.2, the performance is improved by 0.3%. In #4, we also study the effects of bigger training iterations. Increasing the training iterations from 186K to 372K iterations, further improves the performance by 0.2%. All reported results are obtained without ImageNet [52] pretraining.
@@
Correlation Analysis. We sampled 10 subnetworks, training each from scratch for 40,000 iterations. The results, shown in Table I, reveal a high correlation between the mIoU scores within the supernetwork and after retraining, with a Kendall’s $\tau$ of 0.58 and a Pearson Correlation Coefficient ($\rho$) of 0.79. This strong correlation confirms our ability to identify high-performing subnetworks within the supernetwork. Moreover, our search method not only reduces search and GPU memory costs compared to state-of-the-art methods but also efficiently yields multiple viable architectures from a single search process.
V Conclusion

### v2

@@ 3.4 Searching with Multi-objective
Inspired by [29], we formulate the search for the optimal architecture $\alpha^{*}$ as a two-stage optimization problem. The architecture search space $\mathcal{A}$ is encoded in the supernet, denoted as $\mathcal{N}(\mathcal{A},\mathcal{W})$. Each candidate architecture $\alpha\in\mathcal{A}$ shares its weights across $\mathcal{W}$. In the first stage, we optimize the supernet weights $\mathcal{W}$ by:
$\mathcal{W}=\mathop{\mathrm{arg\,min}}\limits_{\mathcal{W}}\mathcal{L}_{train}(\mathcal{N}(\mathcal{A},\mathcal{W}))$ (1)
@@
The prior distribution of $\Gamma(\mathcal{A})$ is important. In order to meet our goal of searching for architectures with arbitrary branches from different resolutions, we pre-define this prior in the training process of the supernet. Since the sub-network with more branches contains the sub-network with fewer branches, we empirically set the sub-network with more branches to have a higher sampling probability.
$\rho_{i}<\rho_{i+1},\quad i\in\{1,2,3\}$ (2)
where $\rho_{i}$ means the sampling probability of architectures containing $i$ branches. Other components in the sub-network $\alpha$ are sampled from a uniform distribution.
Once the supernet training is finished, the second stage is to search for architectures by ranking the accuracy and latency of each sub-network $\alpha\in\mathcal{A}$. We convert the architecture searching problem to a standard multi-objective optimization (MOP) problem.
@@
Where $\alpha$ represents the decision vector, i.e., the searched network architecture, $y$ represents the target vector, i.e., the mIoU and latency. $\mathcal{A}$ represents the search space formed by the decision vector $\alpha$, and $Y$ represents the target space formed by the target vector $y$. The constraint $e(x)\leq 0$ determines the feasible value range of the decision vector. In order to ensure that the searched architectures are effective, we set the following four constraints: 1) The number of paths in each sub-network needs to be greater than 0; 2) For a sub-network that contains multiple paths, the locations of down-sampling should be different; 3) In the down-sampling position, skip connections cannot be selected; 4) The memory usage of a sub-network during training cannot exceed the hardware limit.
Different from the single-objective optimization problem, there is usually no unique optimal solution in an MOP problem, but a Pareto optimal solution set. Hence, we adapt the popular evolutionary algorithm NSGA-II [30] to solve the optimization problem. The details of our search algorithm can be found in the supplemental material.
The detailed search algorithm is presented in Algorithm 1. One-hot encoding is employed to encode the sub-network $\alpha$. When the sampling is over, we will decode the sub-network to evaluate its corresponding mIoU and latency. We solve this multi-objective optimization (MOP) problem under the NSGA-II [30] framework.
@@
4 Experiments
We evaluate our method in both semantic segmentation task and the panoptic segmentation tasks. Panoptic segmentation was introduced by [31] to combine semantic and instance segmentation, in which all object instances are uniquely segmented to make the task even more challenging.
Figure 3: The searched architectures under different objectives on the Cityscapes validation set. Our proposed HyCTAS approach can search for multiple efficient architectures with arbitrary branches in a single run. The color from light to dark indicates that our search results are getting better and better as the search progresses. The solid line represents architectures on the Pareto frontier. Best viewed in color.
Figure 4: An illustration of our searched architectures with multiple branches.
4.1 Benchmark and Evaluation Metrics
@@
Cityscapes [32] is a popular dataset that contains a diverse set of stereo video sequences recorded in street scenes from 50 different cities. It has 5,000 high-quality pixel-level annotations. There are 2,975 images for training, 500 images for validation, and 1,525 images without ground truth for testing to ensure a fair comparison. The dense annotation contains 19 classes for each pixel.
COCO [33] is a much larger and more challenging dataset for object detection and segmentation. There are a total of 133 classes, which include 80 ’thing’ and 53 ’stuff’ classes. We train our models on the train2017 set (118k images) and evaluate on the val2017 set (5k images).
ADE20K [34] has 20k images for training, 2k images for validation, and 3k images for testing. It is used in the ImageNet scene parsing challenge 2016 and has 150 classes and diverse scenes with 1,038 image-level annotations.
@@
Evaluation Metrics. The mean intersection over union per class (mIoU), panoptic quality (PQ), and FPS (frames per second) are used as the metrics for semantic segmentation, panoptic segmentation, and speed, respectively.
4.2 Setup
@@
Search. During training of the supernet, the sampling ratios for $\rho_{1}$, $\rho_{2}$, and $\rho_{3}$ are set to 0.2, 0.3, and 0.5, respectively. The initial channel count for the supernet is 64, and the mini-batch size is set to 16. The supernet is trained on the Cityscapes dataset with $256\times 512$ images from the training set. We use the SGD optimizer with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$ to optimize the supernet weights. The learning rate is set to 0.02 with an exponential learning rate decay of 0.992. The entire supernet optimization takes about 120 hours on a single V100 GPU.
In the search process, the initial population $P_{0}$ is set to 40 and generated by uniform random sampling. The probabilities of crossover and mutation operations are set at 0.9 and 0.02, respectively. Before the inference of a sub-network, the statistics of all Batch Normalization operations are recalculated on a random subset of training data (1000 images in Cityscapes). The number of generations is 20, and we search sub-networks with different branch numbers separately. Hence, a total of 2400 network sub-networks are searched.
@@
Retrain. Similar to [6], we also use the filter multiplier to control the model capacity, and the default values for HyCTAS-S and HyCTAS-L are set to 12 and 16, respectively. The final output stride of HyCTAS-S is 8, and the other big models’ final output strides are set to 4 by merging the low-level features in the stem. We use the SGD optimizer with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$ to optimize the searched networks. The ’poly’ learning rate policy is employed with an initial learning rate of 0.005. Data augmentations on all datasets, i.e., random crop, flipping, etc., are consistent with [35]. On the Cityscapes dataset, the training images are cropped to $512\times 1024$ and the mini-batch size is set to 32. It is worth noting that all our results are trained from scratch without ImageNet pretraining, and we only use the training set without any extra data.
@@
Evaluation. We use an Nvidia GeForce RTX 3090 and an Intel Xeon E5-2680-v4 CPU (2.40GHz) for benchmarking the computation cost. When testing the inference speed, an image with a batch size of 1 is first loaded into the GPU memory and the model is warmed up to reach a steady speed. After that, the inference time is measured by running the model for six seconds. All the reported results, unless specified, are evaluated without multi-scale inference and flipping. The mean intersection over union per class (mIoU), panoptic quality (PQ) are used as the evaluation metrics for semantic segmentation and panoptic segmentation, respectively.
@@
Table 3: Comparison of semantic segmentation on the Cityscapes. $\dagger$: speed on NVIDIA Xavier, $\sharp$: speed with TensorRT acceleration, $\star$: trained from scratch.
@@
$\text{Dynamic}_{L32}$ [24] 1024x2048 79.2 80.0 242.3 - -
$\text{SETR}_{PUP}$ [12] 768x768 79.3 - 318.3 - -
[…]
$\text{OCR}_{HRNetV2-W48}$ [28] 1024x2048 80.6 - 617.2 63.8 4.6
[…]
4.3 Model Exploration
We begin our experiments with the exploration of our searched architectures. As mentioned in the previous section, our search framework is capable of optimizing towards multiple different objectives. To better understand the architecture distributions under different objectives, we conduct three different searches, which cover the objectives ”FLOPs vs. mIoU,” ”Parameters vs. mIoU,” and ”Speed vs. mIoU,” respectively. As shown in Figure 3, we visualize the evolution processes for different objectives. As our method is capable of searching sub-networks with an arbitrary number of branches, we visualize the group of architectures with different numbers of branches in different colors. We also plot the Pareto boundary at the end of the search. It is interesting to find out that searching for different objectives largely affects the architecture distributions. For the following experiments, we choose architectures on the Pareto boundary from searching under ”Speed vs. mIoU” because of better model separation and closer alignment to real-world requirements.
We also visualize the searched architectures with multiple branches in Figure 4. It is interesting to observe how the self-attention module is placed under a different number of branches. When the number of branches is 1, the self-attention modules can be placed in higher resolution to increase the network’s performance; when the number of branches is 3, the self-attention modules would be placed at a deeper stage to better fuse the features from multiple branches. This phenomenon well aligns with self-attention’s advantages and limitations.
4.4 Comparison to SOTA Models
Table 4: Comparisons of panoptic segmentation on Cityscapes validation set. $\dagger$: speed on Tesla V100-SXM2 GPU, $\ddagger$: test with multi-scale and flipping augmentation, $\star$: trained from scratch.
Methods Input Size PQ AP mIoU
$\text{TASCNet}^{\ddagger}$ [38] - 60.4 39.1 78.7
Panoptic FPN [31] - 58.1 33.0 75.7
AUNet [39] - 59.0 34.4 75.6
UPSNet [40] 1024x2048 59.3 33.3 75.2
$\text{UPSNet}^{\ddagger}$ [40] - 60.1 33.3 76.8
Seamless [41] 1025x2049 60.3 33.6 77.5
AdaptIS [42] 1024x2048 62.0 36.3 79.2
DeeperLab [43] 1025x2049 56.5 - -
$\text{SSAP}^{\ddagger}$ [44] - 61.1 37.3 -
[…]
@@ 4.4 Comparison to SOTA Models
Cityscapes. As shown in Table 3, we can see that when the speed is more than 100 FPS, our model establishes new state-of-the-arts. HyCTAS-M consistently outperforms recent BiSeNet [16], Fast-SCNN [17], and DF1-Seg-d8 [36]. In particular, our HyCTAS-M achieves 77.1% mIoU on the Cityscapes validation set with 29.4G FLOPs and 2.4M parameters, surpassing HRNetV2-W18-Small-v1 by 6.8% with a similar model size. On the Cityscapes test set, our HyCTAS-M surpasses FasterSeg [7] by 4.2%. Moreover, the speed of our model reaches an astonishing 87.8 FPS without TensorRT acceleration. Meanwhile, our HyCTAS-S has the smallest FLOPs and parameters and achieves the fastest inference speed at 159.2 FPS. This proves that our models are very competitive when deployed on real hardware.
Without any pretraining or knowledge distillation, our big model HyCTAS-L achieves 80.9 mIoU on the Cityscapes validation set, which is 0.6% higher than Auto-DeepLab [6], while using fewer FLOPs. Compared to the expert-designed HRNetV2-W40 [5], HyCTAS-L is still 0.7% ahead on the validation set with nearly 100G fewer FLOPs. Notably, the HyCTAS-L can still run at 12.4 FPS. Since we target fast semantic segmentation, we did not use dilated separable convolution for fast inference speed. Besides, the mIoU is measured without any evaluation augmentation like flipping or multi-scale.
Compared with recent NAS-related models, our method also has obvious advantages in search time and GPU memory overhead: 1) Thanks to our multi-objective searching strategy, our HyCTAS only takes 5 GPU days to search a series of models which can satisfy different demands, i.e., low latency model or high-performance model. By contrast, other NAS methods can only search for one model at a time, so their search overhead will increase linearly as the number of searched models increases. 2) Our model also achieves better performance than the baseline NAS methods. For example, HyCTAS-M is 3.8% superior to FasterSeg [7] on the Cityscapes validation set.
Compared with the latest development of Transformer-like models [12, 28], our HyCTAS-L model can afford inference using a much larger image and requires fewer FLOPs respectively. This demonstrates the overall advantages of our architectures, which can better combine convolution and self-attention to maximize efficiency.
We also evaluate HyCTAS-L on the Cityscapes validation set with the original image resolution of $1024\times 2048$. In Table  4, we see the superior PQ and speed of our HyCTAS-L. Without any inference tricks, our HyCTAS-L achieves 63.3% PQ and 81.3% mIoU, which is comparable to Panoptic-Deeplab with the HRNet-W48 [5] backbone. However, our HyCTAS-L runs twice as fast as Panoptic-Deeplab with the HRNet-W48 backbone. When we use inference data augmentation, our method achieved the best results. Specifically, our HyCTAS-L surpasses Panoptic-Deeplab [35] by 0.7% on PQ and 0.3% on mIoU. These results are achieved with only Cityscapes fine-annotated images, without using any extra data (coarse-annotated images, COCO, ImageNet, etc.).
Table 5: Comparison of semantic segmentation on the ADE20K.
Methods Encoder FLOPs Params Speed mIoU
(G) (M) (FPS) (%)
FCN [13] MobileNetV2 39.6 9.8 64.4 19.7
PSPNet [45] MobileNetV2 52.9 13.7 57.7 29.6
DeepLabV3+ [46] MobileNetV2 69.4 15.4 43.1 34.0
SegFormer [47] MiT-B0 8.4 3.8 50.5 37.4
FCN [13] ResNet-101 275.7 68.6 14.8 41.4
EncNet [48] ResNet-101 218.8 55.1 14.9 44.7
PSPNet [45] ResNet-101 256.4 68.1 15.3 44.4
DeepLabV3+ [46] ResNet-101 255.1 62.7 14.1 44.1
Auto-DeepLab [6] NAS-F48-ASPP - - - 44.0
HyCTAS(Ours) HyCTAS-L 72.7 43.2 71.8 44.9
ADE20K. We further evaluate our HyCTAS-L on the ADE20K validation set with the original image resolution of $640\times 640$. From Table 5, we can see that our model is much smaller and faster than others. Specifically, our HyCTAS-L achieves 44.9% mIoU with 71.8 FPS, which is 0.9% better than Auto-DeepLab [6]. Compared to those models with a ResNet-101 [49] encoder, we also achieve leading performance while using less than half the FLOPs. For example, the FLOPs of our HyCTAS-L is 72.7G, which is less than half of that of DeepLabV3+ [46]. For the MobileNetV2 [50] encoder, we have similar FLOPs, but our HyCTAS-L surpasses DeepLabV3+ [46], PSPNet [45], and FCN [13] by more than 10%. Furthermore, our HyCTAS-L surpasses the latest SegFormer [47] by 7.3% while also running around 40% faster.
COCO. We employ the recent bottom-up panoptic segmentation framework [35] as our baseline. To ensure fast running speed, we remove the heavy context module and decoder module in Panoptic-Deeplab, which are built based on Atrous Spatial Pyramid Pooling (ASPP) [14]. We further transfer the discovered architecture, i.e., HyCTAS-L, from Cityscapes to train on COCO. To ensure the inference speed, we set the test size to $640\times 640$. Table  6 reveals that without sacrificing much accuracy, our HyCTAS-L achieves an FPS of 63.9. This high speed is over 90% faster than the closest competitor in FPS [43]. Meanwhile, we surpass Deeperlab-MNV2 [43] by a large margin, i.e., 7.5% PQ. Compared to Panoptic-Deeplab-MNV3 [35], our HyCTAS-L surpasses it on multiple performance points with 5.9% PQ. When using multi-scale and flipping, we achieve the best PQ at $640\times 640$ resolution, which is 1.6% better than Panoptic-Deeplab-ResNet-50 and 3.0% better than Deeperlab-Xception-71  [43].
Table 6: Comparison of panoptic segmentation on the COCO val2017 set. $\dagger$ denotes the speed on Tesla V100-SXM2 GPU, and $\ddagger$ denotes test with multi-scale and flipping augmentation.
Methods Input Size Speed PQ $\text{PQ}^{\text{Th}}$ $\text{PQ}^{\text{St}}$
(FPS) (%) (%) (%)
$\text{Deeperlab}_{MNV2}$ [43] 641x641 $17.2^{\dagger}$ 27.9 - -
$\text{Deeperlab}_{Xception-71}$ [43] 641x641 $10.6^{\dagger}$ 33.8 - -
[…]
HyCTAS-L 640x640 63.9 35.4 36.8 33.1
[…]
4.5 Ablation
@@
Compared to Random Sampling. To verify the effectiveness of our searching strategy, we compared our discovered architectures with randomly sampled architectures on the Cityscapes validation set. We randomly sampled 5 architectures with 1-3 branches each from the search space. The results are presented in Table 7, in which each value is the average result of 5 random architectures. The discovered one-branch architecture, Searched-1, outperforms the random one-branch architecture by 4.3% with a smaller model size. The discovered Searched-2 and Searched-3 both surpass random architectures by more than 3%.
@@
Robustness of Searching. We further conducted an ablation study by intentionally changing components to verify the searching robustness of our HyCTAS-M model. As shown in Table 8, we distorted the searched architectures by replacing all self-attention with convolution, replacing all convolution with self-attention, shifting self-attention to another position, and shifting self-attention to another branch, i.e., moving it to a random position on the same stride. It can be seen that our searched models consistently perform better than these variants, proving their robustness.
@@
Table 7: Ablation study on the randomly sampled architectures on the Cityscapes validation set.
@@
Table 8: Ablation study on the searched architectures on the Cityscapes validation set.
@@
Table 9: Experiments with different settings on the Cityscapes validation set. ‘SDP’ denotes Scheduled Drop Path.
@@
Retrain Settings. From Table 9, we can see that adding the scheduled drop path [6] improves the mIoU by 1.5%. When the scheduled drop path probability is set to 0.2, the performance is improved by 0.3%. In #4, we also study the effects of larger training iterations. Increasing the training iterations from 186K to 372K further improves the performance by 0.2%. All reported results are obtained without ImageNet [51] pretraining.
@@
Correlation Analysis. We sampled 10 subnetworks, training each from scratch for 40,000 iterations. The results, shown in Table 1, reveal a high correlation between the mIoU scores within the supernetwork and after retraining, with a Kendall’s $\tau$ of 0.58 and a Pearson Correlation Coefficient ($\rho$) of 0.79. This strong correlation confirms our ability to identify high-performing subnetworks within the supernetwork. Moreover, our search method not only reduces search and GPU memory costs compared to state-of-the-art methods but also efficiently yields multiple viable architectures from a single search process.
Figure 5: Visualization on COCO and Cityscapes validation set. The first row corresponds to COCO original images and the second row presents their semantic segmentation results. The third row corresponds to Cityscapes original images and the fourth row presents their semantic segmentation results. Best viewed in color.
4.6 Visualization
We visualize our predictions in Figure 5. It is obvious to see that our predictions are accurate on fine-grained details thanks to our high resolution representation learning with attention.
5 Limitations
Although our HyCTAS framework achieves state-of-the-art speed and accuracy trade-offs in semantic and panoptic segmentation, several limitations remain:
  • 1.
    Sensitivity to Noise and Occlusions. Our method, like most deep architectures, can degrade in highly noisy or heavily occluded scenarios. Future work could integrate robust loss functions or data augmentation targeted at severe occlusions.
  • 2.
    Extreme Resolution Shifts. While our multi-resolution branches handle moderate scale variations, extremely large input images or drastic changes in scale may still pose memory challenges, even with our memory-efficient attention.
  • 3.
    Domain Generalization. We have mainly evaluated on well-known datasets such as Cityscapes and ADE20K. In real-world applications (e.g., certain industrial or medical imaging domains), domain shifts could undermine performance if the training distribution differs substantially.
We plan to investigate these aspects in future research, potentially by extending our search space to incorporate more robust modules or domain-adaptive strategies.
6 Conclusion